OpenStudy (anonymous):

any1 knows the value of i^i????

OpenStudy (saifoo.khan):

0.207879576 \

OpenStudy (anonymous):

infinity

OpenStudy (anonymous):

how u calculated saifo?

OpenStudy (saifoo.khan):

Calc.

OpenStudy (anonymous):

infinity how?

OpenStudy (anonymous):

sqrt -1^sqrt -1 =infinity

OpenStudy (anonymous):

hm! its i whose root is -1

OpenStudy (anonymous):

its somewhat............................hmmmmmmmm..........................................................................infinity

OpenStudy (anonymous):

its sqrt of (-1) ^ sqrt of (-1)

OpenStudy (anonymous):

infinity???????????how?

OpenStudy (anonymous):

\[i^i\] \[\large {e^{i \frac{\pi}{2} \times i }}\]

OpenStudy (anonymous):

ya thaty i m asking :-)

OpenStudy (anonymous):

u r on d rgt way mgt b ishaan ! thanks

OpenStudy (anonymous):

\[\large{e^{- \frac{\pi}{2}}} \]

OpenStudy (anonymous):

Got your Answer

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

eî pi = -1 -> i pi = ln(-1) ....

OpenStudy (anonymous):

thanks evry1!

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=+i^i